OpenStudy (cloverracer):

John walked 10 miles on Saturday. He walked twice as fast on the second 5 miles of his walk than he walked on the first 5 miles of his walk. Which expression represents the time he spent walking? Let x = John's speed on the first half of his walk.

OpenStudy (cloverracer):

15/x 5/3x 15/2x 10/3x

OpenStudy (cloverracer):

@love10129151

jimthompson5910 (jim_thompson5910):

if x = John's speed on the first half of his walk. , and He walked twice as fast on the second 5 miles of his walk than he walked on the first 5 miles of his walk then what is the speed for the second half of his walk?

OpenStudy (cloverracer):

10..?

jimthompson5910 (jim_thompson5910):

x = speed on first half speed on second half = ??? (he's walking twice as fast as the prev speed)

OpenStudy (cloverracer):

would the answer be 15/2x since the 2 stands for twice and x stands for his speed

jimthompson5910 (jim_thompson5910):

don't worry about the answer choices right now

jimthompson5910 (jim_thompson5910):

2x is the speed on the second half of the walk

jimthompson5910 (jim_thompson5910):

First half D = r*t 5 = x*t 5/x = t t = 5/x ------------- Second half D = r*t 5 = 2x*t 5/2x = t t = 5/2x

jimthompson5910 (jim_thompson5910):

So his time on the first half is 5/x the time on the second half is 5/(2x) add those up to get your answer

OpenStudy (cloverracer):

so it would be 10/3x

jimthompson5910 (jim_thompson5910):

you can't add the denominators like that

jimthompson5910 (jim_thompson5910):

in order to add two fractions, the denominators must be the same and once they are, they don't change

jimthompson5910 (jim_thompson5910):

example